When potassium dichromate (K2CrO4) is added to a solution containing 0.500g silver nitrate (AgNO3) solid silver chromate (Ag2CrO4) is formed


1
Expert's answer
2022-02-18T09:35:02-0500

The balanced equation is below:

K2CrO4(aq) + 2AgNO3(aq) --> Ag2CrO4(s) + 2KNO3(aq)


It is unclear what is required in this problem, but the mass of AgNO3 allows to determine the mass of the precipitate mentioned in the question.

Molar masses are:

AgNO3 - 169.9 g/mol

Ag2CrO4 - 331.7 g/mol


"m(Ag_2CrO_4)=0.500\ g(AgNO_3)\times\frac{1\ mol(AgNO_3)}{169.9\ g(AgNO_3)}\times\frac{1\ mol(Ag_2CrO_4)}{2\ mol(AgNO_3)}\times\frac{331.7\ g(Ag_2CrO_4)}{1\ mol(Ag_2CrO_4)}=0.488\ g"